OpenStudy (anonymous):

ln^3radical33 and ln3=a ln11=b

myininaya (myininaya):

This is hard for me to understand.

myininaya (myininaya):

do you mean \[\ln(\sqrt[3]{33}) \\ \ln(3)=a \\ \ln(11)=b \]? what do you want to do with this exactly though?

myininaya (myininaya):

do you want to write the first line in terms of a and b?

myininaya (myininaya):

if so recall that \[\sqrt[a]{x^b}=x^\frac{b}{a}\]

myininaya (myininaya):

then use power rule for log

myininaya (myininaya):

then use product rule for log

OpenStudy (anonymous):

sorry this is what its asking me. Suppose that ln3=a and ln11=b. use properties of logarithms to write the logarithm in terms of a and b. ln^3radical33

myininaya (myininaya):

\[\sqrt[3]{33}=33^\frac{1}{3} \\ \text{ so you have } \\ \ln(33^\frac{1}{3})\] and what I'm saying is you can use power rule there

myininaya (myininaya):

bring down the exponent in front that is power rule ln(x^r)=rln(x)

myininaya (myininaya):

\[\ln([f \cdot g]^r)=r \ln(f \cdot g) \text{ by power rule } \\ =r [\ln(f)+\ln(g)] \text{ by product rule }\]

myininaya (myininaya):

these are exactly the steps you need to take

myininaya (myininaya):

realize that you can use 33=3*11 somewhere

OpenStudy (anonymous):

so the answer would be 1/2ln(33^1/3)??

OpenStudy (anonymous):

I still need help. i still dont get it